OpenStudy (firejay5):

What are the 4 vertices of the box of this hyperbola? Give the coordinate points.

OpenStudy (firejay5):

\[\frac{ (x+5)^2 }{ 9 } - \frac{ (y+2)^2 }{ 4 } = 1\]

OpenStudy (firejay5):

@satellite73

OpenStudy (firejay5):

I already know two points of the dotted box

OpenStudy (firejay5):

(-8, -2) & (-2, -2)

OpenStudy (anonymous):

yeah i just got to that point myself

OpenStudy (firejay5):

I need the other 2 points so I can draw my dotted box

OpenStudy (anonymous):

you need 4 points

OpenStudy (firejay5):

yes

OpenStudy (firejay5):

the equation of asymptote is: y = -2 +/- 2 over 3 (x + 5)

OpenStudy (anonymous):